11% of the population is left-handed, find from a SRS of n=8
p(exactly 1 is left handed)
P(at most 1 is left-handed)
P(at least 2 are left handed)

Answer by stanbon(75887)

11% of the population is left-handed, find from a SRS of n=8
p(exactly 1 is left handed) = 8C1*0.11*0.89^7 = binompdf(8,0.11,1) = 0.3892
P(at most 1 is left-handed) = binomcdf(8,0.11,1) = 0.7829
P(at least 2 are left handed) = 1 - 0.7829 = 0.2171
----
Cheers,
Stan H.
-----------